# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 6.9 Boston Mechanism
One of the biggest lessons from my experience with the Turkish authorities was this: the availability of a good mechanism, even if it is theoretically perfect, is still not enough by itself to convince policymakers to undertake a costly reform. One must also demonstrate that the mechanism they currently use is clearly inferior to the proposed alternatives, based on criteria that matter to the stakeholders.
@@
In my earlier interactions with the Turkish authorities, I simply failed to establish sufficient value for replacing MCSD with DA. At least with the 1997 data, the outcome of MCSD was the same as that of DA. In theory, this equivalence could have been an artifact of potential preference manipulation by students, but there was no apparent reason—such as field evidence—to believe this was the case. Maybe, I thought, MCSD wasn’t too bad after all… So, I shifted my attention to finding real-life mechanisms that were more visibly flawed according to criteria likely to be important to the stakeholders.
@@
Among the mechanisms we documented in Abdulkadiroğlu and Sönmez (2003) across more than a dozen major U.S. school districts, one that stood out was the mechanism used to allocate K-12 public school seats at Boston Public Schools (BPS). In fact, it was the only mechanism we identified at the time that had been implemented in multiple districts. After researching its history, we concluded that Boston was likely the first district to adopt it. As a result, we named it the Boston mechanism. [^55] Today, the Boston mechanism is also referred to as the immediate acceptance mechanism, reflecting its algorithmic dynamics and contrasting with Gale and Shapley’s renowned Deferred Acceptance algorithm. Other major school districts which used this mechanism included Cambridge, Charlotte, Denver, Minneapolis, Seattle, [^56] See Section 13.2 for Seattle’s unusual experience with Boston and DA mechanisms. and St. Petersburg-Tampa.
@@
In terms of its mechanics, the Boston mechanism is one of the simplest direct mechanisms one can imagine for school choice. Under this procedure the outcome of the mechanism is determined in several steps, first by assigning seats at the first-choice schools of students, next by assigning students unmatched from the first step seats at their second-choice schools, and so on. At every step of the procedure, any given school admits the highest priority students in consideration until all its seats are awarded. Informally speaking, the mechanism first handles the first choices, then the second choices, etc., literally in the most obvious way.
@@
To its credit, if we take the submitted student preferences at face value—that is, assuming they report their preferences truthfully—the outcome of the Boston mechanism is Pareto efficient. After all, this mechanism directly prioritizes accommodating students’ first choices over their second choices, and similarly, it gives higher priority to any given ranked choice compared to those ranked lower in their submitted preferences.
@@
However, at the same time, the Boston mechanism creates strong incentives for students to manipulate their preferences due to these same mechanics. Since it prioritizes assigning seats at any school to students who rank it first in their submitted preferences, it is often unwise for students to “waste” their first choice on schools they are unlikely to secure due to their low priority at those target schools.
@@
We soon discovered that we were not the first to make this observation. For example, Glazerman and Meyer (1994) discussed this issue in the education literature:
@@
Indeed, some school districts using the Boston mechanism explicitly advised parents to strategize when submitting their preferences. As reported in Abdulkadi̇roğlu et al. (2006), the 2004 Boston Public Schools guide states:
@@
These discoveries convinced me that school districts using the Boston mechanism might be more receptive to our reform ideas than the Turkish authorities had been. However, before proceeding, I decided to build a much stronger case against the Boston mechanism than I had been able to make against the MCSD. Specifically, I decided to explore how the Boston mechanism’s vulnerability to strategic manipulation affects its efficiency.
Since the Boston mechanism is not strategy-proof, analyzing its efficiency is not a straightforward exercise. In a laboratory experiment Chen and Sönmez (2006), we examined both the efficiency and the extent to which students engaged in strategic behavior for the Boston mechanism and its competitors DA and TTC. Our experimental analysis revealed that, while the Boston mechanism is Pareto efficient in the absence of strategic manipulation, in our designed environment, it has lower efficiency than both DA and TTC. Indeed, in our experiment, DA outperformed not only the Boston mechanism but also TTC. [^57] While DA and TTC are both strategy-proof, the subjects were not given this important information in our experiment. As a result, many of them engaged in strategic behavior under these mechanisms and reduced their efficiency, but more so for TTC than DA. It is important to emphasize that, in real-life implementation, it is in the best interest of a school district to educate participants on the strategy-proofness of these mechanisms. Indeed, one of the most important benefits of adopting a strategy-proof mechanism is that a school district can encourage students to be truthful without taking any risk. Hence, by not educating the subjects on the strategy-proofness of DA and TTC, in our experimental setup, we analyzed these mechanisms in an environment that is unfavorable for them. Despite this treatment, both mechanisms outperformed the Boston mechanism, thus increasing our confidence that they are likely to be more efficient than the Boston mechanism in real-life implementation. There is, however, disagreement on this point in some of the follow-up literature. For example, using approaches from Bayesian mechanism design, Abdulkadiroğlu et al. (2011) argue that the Boston mechanism can be more efficient than DA in some environments. In order to further support the experimental evidence in Chen and Sönmez (2006), in Ergin and Sönmez (2006), we also showed that DA outperforms the Boston mechanism in terms of efficiency with a complete information equilibrium analysis of the latter mechanism.
To further support the experimental evidence in Chen and Sönmez (2006), we demonstrated in Ergin and Sönmez (2006) that DA outperforms the Boston mechanism in terms of efficiency through a complete information equilibrium analysis of the Boston mechanism. These findings are later formalized in Section 13.1 as Theorem 17 and Corollary 1, where we explore the behavioral aspects of our minimalist market design applications.
@@ 6.10 Policy Impact at Boston Public Schools
With the main analysis in Balinski and Sönmez (1999), Abdulkadiroğlu and Sönmez (2003), the supporting analysis in Chen and Sönmez (2006), Ergin and Sönmez (2006), and anecdotal evidence of strategic behavior, we were ready to make our move against the Boston mechanism. However, it was one additional development that provided us with the impetus to approach the leadership at BPS for a potential reform of their student assignment mechanism.
@@ 6.10.2 Collaboration with Officials at Boston Public Schools
Not only did the Boston Globe story give a significant boost to our policy efforts at BPS, but it also reinforced my belief that I was on the right track with my broader strategy to influence policy. The vulnerability of the Boston mechanism to strategic manipulation was indeed a significant burden for families. Thus, the value of our proposed mechanisms in DA and TTC became more apparent with this story.
@@
During my presentation, it became clear that, more so than anything else, it was the strategy-proofness of DA and TTC that secured this meeting. Given the more clear comparison between DA and the Boston mechanism due to Ergin and Sönmez (2006), in this meeting I focused mostly on DA as a plausible alternative for BPS. The group at BPS, however, were very well-prepared for the papers I had sent, and in the meeting they also inquired about our thoughts on TTC.
@@
In order to support our proposed reform, officials at BPS also provided us with student preference data from earlier years under the Boston mechanism, and requested an empirical analysis of the extent of strategic behavior. This analysis, reported in Abdulkadi̇roğlu et al. (2006), corroborated the presence of strategic behavior under the Boston mechanism, and used by the city to support the reform. [^58] Several subsequent studies, including Pais and Pintér (2008); He (2014); Burgess et al. (2015); Agarwal and Somaini (2018); Dur et al. (2018a); Calsamiglia and Güell (2018); Calsamiglia et al. (2020); Kapor et al. (2020), further validated the prevalence of strategic behavior under the Boston mechanism and its variants. While the task force recommended adoption of the TTC rather than DA, the city at the end decided to adopt DA.
@@
During the public School Committee meeting in May 2005, Boston Public Schools (BPS) officials presented their rationale for their recommendation to abandon the Boston mechanism. This decision was succinctly outlined as follows: [^59] As seen in Figure 3, the details presented here are directly sourced from page 14 of the official 2005 Boston Public Schools presentation to the Boston School Committee.
@@
Consequently, the pivotal factor leading to the abandonment of the Boston mechanism was its susceptibility to strategic manipulation, and the far-reaching consequences of this vulnerability played a decisive role in shaping this course of action. The importance of strategy-proofness in their recommendation was further emphasized as follows: [^60] As seen in Figure 4, the details presented here are directly sourced from page 36 of the official 2005 Boston Public Schools presentation to the Boston School Committee.
@@
The appeal of adopting a strategy-proof mechanism positioned DA and TTC as natural replacements for the Boston mechanism. The officials at BPS explained their recommendation of DA as follows: [^61] As seen in Figure 5, the details presented here are directly sourced from page 37 of the official 2005 Boston Public Schools presentation to the Boston School Committee.
@@
Hence, the desiderata of no justified envy and strategy-proofness were the primary reasons behind the BPS recommendation for DA.
Finally, officials at BPS explained their reluctance to recommend TTC, despite its strategy-proofness and Pareto efficiency, as follows: [^62] As seen in Figure 6, the details presented here are directly sourced from page 38 of the official 2005 Boston Public Schools presentation to the Boston School Committee.
@@
Even when priorities were not determined by standardized exams, officials at BPS were uncomfortable with the idea of trading priorities. As a result, the failure to satisfy no justified envy was seen as a drawback of this mechanism. Additionally, officials were concerned that, while TTC is strategy-proof, this feature might not be clear to families, potentially leading them to engage in strategic behavior.
Officials also expressed concern that the mechanics of TTC were less transparent than those of DA, which are effectively demonstrated through cutoff scores, as established in Theorem 3.
In June 2005, the Boston School Committee voted to replace the Boston mechanism with DA. The city implemented DA starting in the next school year and has used it for the allocation of K-12 public school seats ever since. The critical role played by various axioms in both the successful policy efforts in Boston and the earlier unsuccessful ones in Turkey was instrumental in the initial development of minimalist market design.
@@ 6.11 The 2003 NYC High School Reform: Persuasion Takes a Backseat
The 2005 reform at BPS was one of two high-profile school choice initiatives led by design economists in the mid-2000s in the U.S. The other was a seemingly similar school choice reform in New York City (NYC) for allocating high school seats (Abdulkadiroğlu et al., 2005a). Like Boston, NYC replaced its flawed mechanism with a version of DA in 2003. Similar to the abandoned Boston mechanism, the NYC high school mechanism also suffered from strategic vulnerabilities. Given these parallels, the success of both reforms was attributed to similar factors.
However, while there are clear parallels between these school choice reforms, there are also crucial differences. Most notably, the development of the two reform processes diverged significantly. I argue that the motivations and implementation of the BPS and NYC high school reforms differed fundamentally, despite their surface-level similarities. Unfortunately, much of the literature oversimplifies these cases, focusing heavily on the specifics of the chosen mechanisms while overlooking the deeper dynamics that shaped each reform.
By examining the unique dynamics of each reform, this section highlights the importance of understanding the diverse driving forces behind these initiatives. Such insights are vital for design economists aiming to influence policy through research, as they reveal broader lessons that extend beyond the technical details of the mechanisms themselves.
In 2002, the New York City Department of Education (NYCDOE) faced the following three issues with the assignment of incoming high school students to more than 500 programs (Abdulkadiroğlu et al., 2005a):
@@
    Of nearly 100,000 students, approximately 30% had been assigned to a school that was not included in their submitted preference lists.
@@
    The system was vulnerable to strategic preference manipulation by the students in a way similar to the Boston mechanism.
@@
    A number of schools were able to conceal capacity from the central administration and preserve seats for allocation outside the system. [^63] These phenomena were studied earlier in a two-sided matching setting in Sönmez (1997, 1999a).
Based on these failures, the NYCDOE Director of Strategic Planning reached out to Alvin Roth in May 2003 to explore whether the matching process used for the U.S. medical residency match could be adapted to design a new high school matching process. Similar to the DA mechanism advocated by Balinski and Sönmez (1999) for student placement and by Abdulkadiroğlu and Sönmez (2003) for school choice, the Roth and Peranson design (see Section 3.2.2), used for the U.S. medical residency match, is also based on Gale and Shapley’s celebrated individual-proposing Deferred Acceptance algorithm.
Given the extent of the crisis in NYC and the decision makers’ initial inclination toward the DA mechanism, reinforced by Roth’s advocacy, the city adopted a version of the DA mechanism in October 2003 on a relatively fast track—just a few months after Abdulkadiroğlu and Sönmez (2003) appeared in print—for students starting high school in Fall 2004. As in the BPS school choice reform, the decision makers’ desire to “limit the need to game the system” was clearly highlighted by Peter Kerr, Director of Communications for the NYCDOE, in a New York Times op-ed (Kerr, 2003): [^64] While the strategy-proofness of the DA mechanism was leveraged by the NYCDOE to justify the reform, the city limited the number of schools students could rank to a maximum of 12, thereby adopting a version of DA vulnerable to potential preference manipulation. For an experimental analysis on how such caps on the number of schools that can be ranked affect submitted preferences, see Calsamiglia et al. (2010).
@@
Despite these similarities with the BPS school choice reform, there are notable differences. By the time a team of design economists became involved in the redesign of the high school assignment process in NYC, the need for reform was already well established, and the authorities were seeking expert advice to fix their broken system. Consequently, they enlisted a leading expert in matching market design, with experience in guiding the reform of the U.S. medical match, to redesign NYC’s high school assignment process. Furthermore, the authorities were already inclined to adopt a version of DA.
Because there was no need for an elaborate persuasion strategy to replace New York City’s broken school choice mechanism with DA—only a reinforcement of the officials’ inclinations—the lessons it offers differ from those of the 2005 BPS school choice reform. [^65] The 2003 NYC high school assignment reform is clearly a “commissioned” reform. At the same time, given the similarities in some of the failures between NYC and Boston, as well as the authorities’ pre-existing inclination toward adopting a version of DA, the reform process in NYC closely resembles the process outlined in Abdulkadiroğlu and Sönmez (2003). In this sense, to some extent the 2003 NYC high school assignment reform also lends external validity to Abdulkadiroğlu and Sönmez (2003). Indeed, with the exception of the final tinkering phase, all other steps of the institutional redesign process I outlined in Section 4.2 were bypassed in this reform.
Apart from the practical value it generated for the students in NYC, this reform’s main contributions to the market design literature are twofold: First, it increased the visibility of school choice as a successful application of the field in the early phases of market design. Second, it fostered a rich follow-up empirical school choice literature through papers such as Abdulkadiroğlu et al. (2014) and Abdulkadiroğlu et al. (2017).

### v4

@@ 6.9 Boston Mechanism
One of the biggest takeaways from my experience with the Turkish authorities was that the availability of a good mechanism—even one that is theoretically perfect—is not, by itself, enough to convince policymakers to undertake a costly reform. One must also demonstrate that the mechanism they currently use is clearly inferior to the proposed alternatives, based on criteria that matter to the stakeholders.
@@
In my earlier interactions with the Turkish authorities, I failed to establish sufficient value for replacing MCSD with DA. At least with the 1997 data, the outcome of MCSD was the same as that of DA. In theory, this equivalence could have been an artifact of potential preference manipulation by students, but there was no apparent reason—such as field evidence—to believe this was the case. Maybe, I thought, MCSD wasn’t too bad after all… So, I shifted my attention to finding real-life mechanisms that were more visibly flawed according to criteria likely to be important to the stakeholders.
@@
Among the mechanisms we documented in Abdulkadiroğlu and Sönmez (2003) across more than a dozen major U.S. school districts, one that stood out was the mechanism used to allocate K–12 public school seats at Boston Public Schools (BPS). It was, in fact, the only mechanism we identified at the time that had been implemented in multiple districts. After researching its history, we concluded that Boston was likely the first district to adopt it, and we therefore dubbed it the Boston mechanism. [^64] Today, the Boston mechanism is also referred to as the immediate acceptance mechanism, reflecting its algorithmic dynamics and contrasting with Gale and Shapley’s renowned deferred acceptance algorithm. Other major school districts that used this mechanism included Cambridge, Charlotte, Denver, Minneapolis, Seattle, [^65] See Section 13.2 for Seattle’s unusual experience with the Boston and DA mechanisms. and St. Petersburg–Tampa.
@@
In terms of its mechanics, the Boston mechanism is one of the simplest direct mechanisms one can imagine for school choice. Under this procedure, the outcome is determined in several steps: first, seats are assigned at students’ first-choice schools; next, students who remain unmatched are considered for their second-choice schools, and so on. At each step, every school admits the highest-priority students among those applying until all of its seats are filled. Informally speaking, the mechanism first processes first choices, then second choices, and so forth — literally in the most straightforward way possible.
Figure 13 illustrates the execution of the Boston mechanism for the same setting as in Examples 10 and 12, in which we executed the DA and SC–TTC mechanisms, respectively.
Figure 13: Execution of the Boston mechanism for the setting in Examples 7 and 8. The top-left panel shows the setting. In Step 1, all students propose to their first choices, and all schools admit the top-ranked offers up to capacity (two for $X$, one each for $Y$ and $Z$). Assignments are final. In Step 2, only student B and school $Z$ are active; the rest are shown in faint gray. Since the second choice of B is not in the process, this round becomes obsolete. Student B then proposes to $Z$ as her third choice and is assigned a seat there in Step 3.
@@
To its credit, if we take the submitted student preferences at face value—that is, assuming they report their preferences truthfully—the outcome of the Boston mechanism is Pareto efficient. After all, this mechanism directly prioritizes accommodating students’ first choices over their second choices, and similarly gives higher priority to any given ranked choice compared to those ranked lower in the submitted preferences. Indeed, as can be seen in Figure 13, the outcome of the Boston mechanism coincides with that of SC–TTC—a Pareto-efficient mechanism—for the setting in Example 8.
@@
However, whereas SC–TTC achieves Pareto efficiency as a strategy-proof mechanism, the Boston mechanism creates strong incentives for students to misrepresent their preferences, precisely because of the same simple mechanics that make it Pareto efficient. Since it prioritizes assigning seats at any school to students who rank it first in their submitted preferences, it is often unwise for students to “waste” their first choice on schools they are unlikely to secure due to low priority at those schools. Indeed, in Figure 13, Banu can benefit from misrepresenting her preferences by ranking school $Y$ first—thereby avoiding “wasting” her first choice by ranking her true first-choice $X$ at the top of her preferences under a mechanism that penalizes truthful behavior and ultimately causes her to receive a seat at her last-choice school $Z$.
@@
We soon discovered that we were not the first to observe that Boston mechanism harbored strong incentives for preference manipulation. For example, Glazerman and Meyer (1994) discussed this issue in the education literature:
@@
Indeed, some school districts using the Boston mechanism explicitly advised parents to strategize when submitting their preferences. As reported in Abdulkadiroğlu et al. (2006), the 2004 Boston Public Schools guide states:
@@
These discoveries convinced me that school districts using the Boston mechanism might be more receptive to our reform ideas than the Turkish authorities had been. However, before proceeding, I decided to build a much stronger case against the Boston mechanism than I had been able to make against the MCSD. Specifically, I decided to explore how the Boston mechanism’s vulnerability to preference manipulation affects its efficiency.
Since the Boston mechanism is not strategy-proof, analyzing its efficiency is not a straightforward exercise. In a laboratory experiment Chen and Sönmez (2006), we examined both the efficiency and the extent to which students engaged in strategic behavior for the Boston mechanism and its competitors DA and SC–TTC. Our experimental analysis revealed that, while the Boston mechanism is Pareto efficient in the absence of strategic manipulation, in our designed environment, it has lower efficiency than both DA and SC–TTC. [^66] While DA and SC–TTC are both strategy-proof, the subjects were not given this important information in our experiment. As a result, many of them engaged in strategic behavior under these mechanisms and reduced their efficiency, but more so for SC–TTC than DA. It is important to emphasize that, in real-life implementation, it is in the best interest of a school district to educate participants on the strategy-proofness of these mechanisms. Indeed, one of the most important benefits of adopting a strategy-proof mechanism is that a school district can encourage students to be truthful without taking any risk. Hence, by not educating the subjects on the strategy-proofness of DA and SC–TTC, in our experimental setup, we analyzed these mechanisms in an environment that is unfavorable for them. Despite this treatment, both mechanisms outperformed the Boston mechanism, thus increasing our confidence that they are likely to be more efficient than the Boston mechanism in real-life implementation.
To further support the experimental evidence in Chen and Sönmez (2006), we showed in Ergin and Sönmez (2006) that DA outperforms the Boston mechanism in efficiency, based on a complete-information equilibrium analysis of the latter. These findings are formalized in Section 13.1 as Theorem 16 and Corollary 1, where we explore the behavioral aspects of our minimalist market design applications. [^67] As we elaborate in Footnote 156 in Section 13.1, Ergin and Sönmez (2006) shows that the complete-information assumption is key to the unambiguous efficiency advantage of DA over the Boston mechanism. Indeed, they also show that, in an incomplete-information environment, a student may prefer a Bayesian Nash equilibrium outcome of the Boston mechanism to the truthful dominant-strategy outcome of DA. Along these lines, using tools from Bayesian mechanism design, Abdulkadiroğlu et al. (2011) show that the Boston mechanism can be more efficient than DA in certain environments.
@@ 6.10 Policy Impact at Boston Public Schools
With the main analysis in Balinski and Sönmez (1999) and Abdulkadiroğlu and Sönmez (2003), the supporting analysis in Chen and Sönmez (2006) and Ergin and Sönmez (2006), and anecdotal evidence of strategic behavior, we were ready to make our move against the Boston mechanism (see Table 2 for a summary comparison of Boston, DA, SC–TTC, and MCSD with respect to policy-relevant properties). However, it was one additional development that provided the impetus to approach the leadership at BPS for a potential reform of their student assignment mechanism.
Table 2: Properties of school choice mechanisms.
Property Boston DA SC–TTC MCSD
Strategy-proofness X X
No justified envy X X
Pareto efficiency * § X
Respect for priorities X X
  • Notes:  satisfied; X not satisfied.
  • * Pareto efficiency holds only under truthful preferences (hence vulnerable to preference manipulation).
  • § Pareto efficiency holds in a restricted, second-best sense within the set of outcomes that satisfy no justified envy.
@@ 6.10.2 Collaboration with Officials at Boston Public Schools
Not only did the Boston Globe story give a significant boost to our policy efforts at BPS, but it also reinforced my belief that I was on the right track with my broader strategy to influence policy. The vulnerability of the Boston mechanism to strategic manipulation was indeed a significant burden for families. Thus, the value of our proposed mechanisms in DA and SC–TTC became more apparent with this story.
@@
During my presentation, it became clear that, above all, the vulnerability of the Boston mechanism to preference manipulation, together with the strategy-proofness of its proposed alternatives—DA and SC–TTC—was what had secured this meeting. Given the clearer comparison between DA and the Boston mechanism provided by Ergin and Sönmez (2006), I focused in this meeting mainly on DA as a plausible alternative for BPS. The group at BPS, however, were well prepared for the papers I had sent, and during the meeting they also inquired about our thoughts on SC–TTC.
@@
In order to build a case for reform to the Boston School Committee and other constituents, officials at BPS provided us with student preference data from earlier years under the Boston mechanism and requested an empirical analysis of the extent of strategic behavior. This analysis corroborated the presence of strategic behavior under the Boston mechanism and was used by the city to make the case for reform (Abdulkadiroğlu et al., 2006). [^68] Several subsequent studies, including Pais and Pintér (2008); He (2014); Burgess et al. (2015); Agarwal and Somaini (2018); Dur et al. (2018a); Calsamiglia and Güell (2018); Calsamiglia et al. (2020); Kapor et al. (2020), further validated the prevalence of strategic behavior under the Boston mechanism and its variants. Although the task force recommended adoption of SC–TTC rather than DA, the city ultimately chose to adopt DA.
@@
Figure 14: Selected pages from the Boston Public Schools presentation, Recommendation to Implement a New BPS Assignment Algorithm, presented to the Boston School Committee (May 2005). Panels show (a) page 14, (b) page 36, (c) page 37, and (d) page 38.
During the public School Committee meeting in May 2005, Boston Public Schools (BPS) officials presented their rationale for their recommendation to abandon the Boston mechanism. This decision was succinctly outlined as follows: [^69] All bullet-point lists reproduced in this section are taken verbatim from the official 2005 Boston Public Schools presentation to the Boston School Committee (see Figure 14, Panels (a)–(d)).
@@
Consequently, the pivotal factor leading to the abandonment of the Boston mechanism was its susceptibility to strategic manipulation, and the far-reaching consequences of this vulnerability played a decisive role in shaping this course of action. The importance of strategy-proofness in their recommendation was further emphasized as follows:
@@
The appeal of adopting a strategy-proof mechanism positioned DA and SC–TTC as natural replacements for the Boston mechanism. The officials at BPS explained their recommendation of DA as follows:
@@
Hence, the desiderata of strategy-proofness and no justified envy were the primary reasons behind the BPS recommendation for DA.
Finally, officials at BPS explained their reluctance to recommend SC–TTC, despite its strategy-proofness and Pareto efficiency, as follows:
@@
Even when priorities were not determined by standardized exams, officials at BPS were uncomfortable with the idea of trading priorities. [^70] More specifically, BPS officials were concerned about trading “sibling” priorities, which were intended to allow siblings to share the schooling experience at the same school. Leveraging sibling priorities to secure assignments in other schools was not considered acceptable. In contrast, they were less concerned about trading “neighborhood” priorities. See Rodríguez-Álvarez and Romero-Medina (2024) for an analysis of settings where some priority-awarding student characteristics can be traded while others cannot. As a result, the failure to satisfy no justified envy was seen as a drawback of this mechanism. Additionally, officials were concerned that, while SC–TTC is strategy-proof, this feature might not be clear to families, potentially leading them to engage in strategic behavior.
Officials also expressed concern that the mechanics of SC–TTC were less transparent than those of DA, which are effectively demonstrated through cutoff scores, as established in Theorem 3.
In June 2005, the Boston School Committee voted to replace the Boston mechanism with DA. The city implemented DA starting in the next school year and has used it for the allocation of K–12 public school seats ever since. The critical role played by various axioms in both the successful policy efforts in Boston and the earlier unsuccessful ones in Turkey was instrumental in the initial development of minimalist market design.
@@ 6.11 The 2003 NYC High School Reform: Persuasion Takes a Backseat
The 2005 reform at BPS was one of two high-profile school choice initiatives led by economic designers in the mid-2000s in the U.S. The other was the 2003 New York City (NYC) reform of the flawed high school admissions system (Abdulkadiroğlu et al., 2005a). Both systems had suffered from strategic vulnerabilities, and in each case the reform introduced a variant of DA. Despite these parallels, the processes that produced the two reforms diverged in important ways. Boston’s reform was triggered by academic recommendations, whereas the NYC reform originated within the Department of Education.
In 2002, the New York City Department of Education (NYCDOE) faced three major issues in assigning nearly 100,000 students to more than 500 programs (Abdulkadiroğlu et al., 2005a):
@@
    Approximately 30% of students were assigned to a school not included on their submitted preference lists.
@@
    The system was vulnerable to preference manipulation, similar to the Boston mechanism.
@@
    Some schools concealed capacity from the central administration to reserve seats for allocation outside the system. [^71] These phenomena were studied earlier in a two-sided matching setting in Sönmez (1997, 1999a).
In response to these failures, the NYCDOE Director of Strategic Planning reached out to Alvin Roth in May 2003 to explore whether the matching process used in the U.S. medical residency match could be adapted to high school admissions. Similar to the DA mechanism advocated by Balinski and Sönmez (1999) for student placement and by Abdulkadiroğlu and Sönmez (2003) for school choice, the Roth and Peranson design—developed for the medical residency match (see Section 3.2.2)—is also based on Gale and Shapley’s celebrated individual-proposing deferred acceptance algorithm.
Given the extent of the crisis in NYC and decision-makers’ initial inclination toward DA, reinforced by Roth’s advocacy, the city adopted a version of the mechanism in October 2003 on a fast track—just a few months after Abdulkadiroğlu and Sönmez (2003) appeared in print—for students starting high school in Fall 2004. As in Boston, officials emphasized their desire to “limit the need to game the system,” as highlighted by Peter Kerr, Director of Communications for the NYCDOE, in a New York Times op-ed (Kerr, 2003): [^72] While the strategy-proofness of DA was used to justify the reform, NYC capped the number of schools students could rank at 12, leaving the system vulnerable to strategic behavior. For experimental analysis of such caps, see Calsamiglia et al. (2010).
@@
The critical difference between the school choice reforms in Boston and New York City lies in how they were initiated and carried out. Boston’s reform was driven directly by academic advocacy and followed a thorough due diligence process, whereas NYC decision-makers, already inclined toward DA, commissioned expert guidance to swiftly resolve an ongoing crisis, bypassing such vetting in the process. As such, the NYC case represents both a “commissioned” reform and an instance of external validity for the theoretical framework in Abdulkadiroğlu and Sönmez (2003), given the parallels between the NYC reform and this work. Because swift execution was essential, many stages of the institutional redesign process outlined in Section 4.2 were shortened, handled during implementation, or bypassed altogether.
Beyond helping end a major crisis and delivering tangible benefits for students, the NYC high school admissions reform made two key contributions to the market design literature: first, it increased the visibility of school choice as a successful application of the field in the early phases of market design; second, it catalyzed a rich follow-up empirical literature (e.g., Abdulkadiroğlu et al., 2014, 2017). Together, the commissioned reform in NYC and the research-driven BPS school choice reform highlight two distinct but complementary pathways through which market design research can help shape policy.
While the NYC case validated earlier market design research in school choice, it did so without the persuasion strategies that had been central in Boston or any minimalism apparent in the reform, as much of the abandoned mechanism’s design was not publicly disclosed. Lacking a clear link to two of the framework’s three pillars, it offered little sense of the broader potential of minimalist market design at that stage. This potential became much clearer in the policymaker-driven reforms that followed in England and Chicago, where officials independently identified failures similar to those in Boston, traced them to the same root causes, and arrived at parallel resolutions—notably without any direct involvement from economic designers.